Brilliant, Inc. reported the following results from the sale of 31,000 units of IT-54:

Sales $ 546,000
Variable manufacturing costs 310,000
Fixed manufacturing costs 124,000
Variable selling costs 54,800
Fixed administrative costs 36,400

  
Extra Company has offered to purchase 3,700 IT-54s at $13 each. Brilliant has available capacity, and the president is in favor of accepting the order. She feels it would be profitable because no variable selling costs will be incurred. The plant manager is opposed because the "full cost" of production is $14. Which of the following correctly notes the change in income if the special order is accepted?
Answer #1
Variable manufacturing costs per unit 10 =310000/31000
Incremental revenue 48100 =3700*13
Less: Incremental costs 37000 =3700*10
Net Change in income 11100
$11,100 increase